# Incentive-Compatible Vertiport Reservation in Advanced Air Mobility: An Auction-Based Approach

### v2

@@ 4.3 Reduction to Mixed Binary Linear Program
First, we prove that we can drop (12g) when solving (12). Suppose there exists $r\in R,t\in[H]$, $q\in[\textsf{park}(r,t)-1]$ such that $\textbf{A}^{\dagger}_{q,r,t}<\textbf{A}^{\dagger}_{q+1,r,t}$. By swapping the value of $\textbf{A}^{\dagger}_{q+1,r,t}$ with that of $\textbf{A}^{\dagger}_{q,r,t}$, we get a new feasible allocation with a weakly higher objective value. This is because $\bar{\textbf{w}}_{q+1,r,t}\leq\bar{\textbf{w}}_{q,r,t}$ as $\lambda\geq 0$ and $C_{r,t}(\cdot)$ is discrete convex.
Additionally, we can also relax (12f) to (16c) because, for any feasible value of $\delta$, the optimization problem (12) is an integer linear program where the constraint matrix $\bar{I}_{\star}$ satisfies total unimodularity, so it is guaranteed to have an integral solution (Schrijver, 1998, Chapter 19). ∎
For any fixed values of binary variables $(\delta_{i,j,\tau})_{i\in F,j\in A_{i},\tau\in T_{i,j}^{\textsf{dep}}}$, the optimization problem (12) is in fact a maximum-weight flow problem. Thus, one can enumerate all the departure time combinations, and solve each maximum-weight flow problem with the number of scenarios being $\prod_{i\in F,j\in A_{i}}|\{t^{\textsf{dep}}_{i,j,k}|k\in m_{i,j}\}|$. The complete problem can be solved efficiency using the above MBLP approach, which will provide speed-up due to some techniques implemented in commercial solvers such as branch and bound, cutting-plane methods, etc.
@@ 5 Discussions
We show how the proposed mechanism generalizes existing works in Section 5.1 and present some extensions in Section 5.2.
@@ 5.1 Connections to Existing Mechanisms
Consider $H=1$, $\textsf{arr}(r,1)=\infty$, $\textsf{dep}(r,1)=\infty$, $\forall r\in R$, and $|A_{i}|=1$, $\forall i\in F$.
@@
    Air Traffic Protocol: When we treat each vertiport $r\in R$ as a sector with $\textsf{park}(r,1)$ being the sector capacity, our model generalizes the problem studied in (Qin and Balakrishnan, 2022), where the authors did not consider arrival and departure capacities and assumed fleet operators with single aircraft.
@@ 5.2 Extensions of the Proposed Mechanism
Our framework can be easily extended to incorporate the following aspects.
@@
    External Demand: Aircraft that are not available in the service area of the SP at $t=0$ can be incorporated in our framework by setting $r^{\textsf{orig}}_{i,j}=\mathcal{O}$ and $t^{\textsf{dep}}_{i,j,k}=0,\forall k\in m_{i,j}$ [^8] In this case, $r^{\textsf{orig}}_{i,j}$ and $t^{\textsf{dep}}_{i,j,k}$ do not affect our analysis, so we can set them arbitrarily., where $\mathcal{O}$ denotes an auxiliary vertiport denoting aircraft coming from outside of the service area of SP.
@@
  • (iii)
    Cancellation Policy: It is possible to cancel or re-allocate some of the previously scheduled flights due to changing vertiport capacities or newly emerging aircraft. While there is no single criterion for re-allocation, it is typical to consider three aspects: congestion, efficiency, and fairness, where we cancel flights from congested vertiports, with low valuations, or at random, respectively.
@@ 6 Conclusion
In this work, we propose an auction mechanism to incentivize fleet operators to report their valuations truthfully and consequently perform a socially optimal vertiport reservation. This approach adapts the popular Vickrey–Clarke–Groves mechanism while considering the egalitarian, congestion-aware, and computational issues. The proposed framework could be of interest beyond air traffic management, such as multi-robot coordination. In a follow-up paper, we shall provide a numerical analysis of the mechanism’s performance. Several interesting avenues exist for future research. First, we would like to extend the auction mechanism to include waypoints in airspace, thus moving toward a more complete air traffic flow management formulation. Second, different fairness notions have been considered in airspace and other areas, such as reversals, takeovers, and priority guarantees (Bertsimas and Gupta, 2016; Su et al., 2023). It is interesting to compare different formulations, both theoretically and empirically.
Appendix A Table of Notations

### v3

@@ 4.3 Reduction to Mixed Binary Linear Program
First, we prove that we can drop (12g) when solving (12). Suppose there exists $r\in R,t\in[H]$, $q\in[\textsf{park}(r,t)-1]$ such that $\textbf{A}^{\dagger}_{q,r,t}<\textbf{A}^{\dagger}_{q+1,r,t}$. By swapping the value of $\textbf{A}^{\dagger}_{q+1,r,t}$ with that of $\textbf{A}^{\dagger}_{q,r,t}$, we get a new feasible allocation with a weakly higher objective value. This is because $\bar{\textbf{w}}_{q+1,r,t}\leq\bar{\textbf{w}}_{q,r,t}$ as $\lambda\geq 0$ and $C_{r,t}(\cdot)$ is discrete convex. Then, for any feasible value of $\delta$, the optimization problem (12) is an integer linear program where the constraint matrix $\bar{I}_{\star}$ satisfies total unimodularity, so it is guaranteed to have an integral solution (Schrijver, 1998, Chapter 19). ∎
For any fixed values of binary variables $(\delta_{i,j,\tau})_{i\in F,j\in A_{i},\tau\in T_{i,j}^{\textsf{dep}}}$, the optimization problem (12) is a maximum-weight flow problem. Thus, one can enumerate all the departure time combinations, and solve each maximum-weight flow problem with the number of scenarios being $\prod_{i\in F,j\in A_{i}}|\{t^{\textsf{dep}}_{i,j,k}|k\in m_{i,j}\}|$. The complete problem can be solved efficiency using the above MBLP approach, which will provide speed-up due to some techniques implemented in commercial solvers such as branch and bound, cutting-plane methods, etc.
@@ 5 Discussions
We show how the proposed mechanism generalizes existing works in Subsection 5.1 and present some extensions in Subsection 5.2.
@@ 5.1 Connections to Existing Mechanisms
We consider $H=1$, $\textsf{arr}(r,1)=\infty$, $\textsf{dep}(r,1)=\infty$, $\forall r\in R$, and $|A_{i}|=1$, $\forall i\in F$.
@@
    Air Traffic Protocol: When we treat each vertiport $r\in R$ as a sector with $\textsf{park}(r,1)$ being the sector capacity, our model generalizes the problem studied in (Qin and Balakrishnan, 2022), where the authors did not consider arrival and departure capacities and assumed single-aircraft FOs.
@@ 5.2 Extensions of the Proposed Mechanism
    Arrival, Departure, and Airborne Congestion: To consider congestion due to arriving and departing aircraft, we can apply the same technique in $\bar{E}_{3}$ to $\bar{E}_{1}$ and $\bar{E}_{2}$ by constructing corresponding edge weights. To consider airborne congestion, we treat waypoints in the airspace as vertiports and setting corresponding capacities and congestion costs.
@@
    External Demand: Aircraft that are not available in the service area of the SP at $t=0$ can be incorporated in our framework by setting $r^{\textsf{orig}}_{i,j}=\mathcal{O}$ and $t^{\textsf{dep}}_{i,j,k}=0,\forall k\in m_{i,j}$ [^11] In this case, $r^{\textsf{orig}}_{i,j}$ and $t^{\textsf{dep}}_{i,j,k}$ do not affect our analysis, so we can set them arbitrarily..
  • (iii)
@@
  • (iv)
    Cancellation Policy: It is possible to cancel or re-allocate some of the previously scheduled flights due to changing vertiport capacities or newly emerging aircraft. While there is no single re-allocation policy, it is typical to consider three aspects: congestion, efficiency, and fairness, where we cancel flights from congested vertiports, with low valuations, or at random, respectively.
@@ 6 Conclusion
In this work, we propose an auction mechanism to incentivize fleet operators to report their valuations truthfully and consequently perform a socially optimal allocation of vertiport access. This approach adapts the popular Vickrey–Clarke–Groves mechanism while considering the egalitarian, congestion-aware, and computational issues. The proposed framework could be of interest beyond air traffic management, such as multi-robot coordination. In a follow-up paper, we shall provide a numerical analysis of the mechanism’s performance. Several intriguing avenues exist for future research. First, we would like to extend the auction mechanism to include waypoints in airspace, thus moving toward a more complete air traffic flow management formulation. Second, a careful analysis of the effect of flight operator weights in our proportional fairness metric is needed. Also, different fairness notions have been considered in airspace and other areas, such as reversals, takeovers, and priority guarantees (Bertsimas and Gupta, 2016; Su et al., 2023). It is interesting to compare different formulations, both theoretically and empirically.
Appendix A Proof of Lemma 4.2
Proof.
First, note that any feasible solution to (12b)-(12g) has the same value of $\textbf{A}_{e}(x)$ for $e\in\bar{E}_{6}\cup\bar{E}_{7}\cup\bar{E}_{9}$ since the lower and upper bound on capacity are the same on these edges by construction. Thus, it is sufficient to show that the values of $\textbf{A}_{e}$ for $e\in\bar{E}_{3}\cup\bar{E}_{5}\cup\bar{E}_{6}\cup\bar{E}_{7}\cup\bar{E}_{8}\cup\bar{E}_{9}$ uniquely determine a feasible solution $(\textbf{A},\delta)$ that satisfies (12b)-(12g). Particularly, we will show that we can uniquely recover the values of $\textbf{A}_{e}$ for $e\in\bar{E}_{1}\cup\bar{E}_{2}\cup\bar{E}_{4}.$
Vertex Incoming Edges Outgoing Edges
$\bar{\nu}^{\textsf{arr}}(r,t)$ ${\bar{E}_{5}}$ $\bar{E}_{1}$
$\bar{\nu}(i,j,\tau)$ $\bar{E}_{4}$ ${\bar{E}_{5}}$
$\bar{\nu}^{\textsf{dep}}(r,t)$ $\bar{E}_{2}$ ${\bar{E}_{4}}$
$\bar{\nu}(i,j,0)$ $\bar{E}_{7}$ $\bar{E}_{9}$
Note that flow balance at nodes of the form $\bar{\nu}^{\textsf{arr}}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{1}$, as we know these values for edges $\bar{E}_{5}$. Next, flow balance at nodes of the form $\bar{\nu}(i,j,\tau)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{4}$, as we know these values for edges $\bar{E}_{5}$. This and the capacity constraints on $\bar{E}_{4}$, ensure that we know the value of $\delta$. Next, flow balance at nodes of the form $\bar{\nu}^{\textsf{dep}}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{2}$, as we can uniquely determine these values on $\bar{E}_{4}$. Finally, flow balance at nodes of the form $\bar{\nu}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{2}$, as we can uniquely determine these values on $\bar{E}_{1}\cup\bar{E}_{3}\cup\bar{E}_{6}\cup\bar{E}_{8}\cup\bar{E}_{9}$.
Appendix B Table of Notations
@@
ACKNOWLEDGMENTS
Chinmay Maheshwari, Pan-Yang Su and Shankar Sastry acknowledge support from NSF Collaborative Research: Transferable, Hierarchical, Expressive, Optimal, Robust, Interpretable NETworks (THEORINET) Award No. DMS 2031899. We thank Maria Gabriela Mendoza and Hamsa Balakrishnan for the helpful discussions.